Malibu Corporation has monthly fixed costs of $62,000. It sells two products for which it has provided the following information.

Sales Price Contribution
Margin
Product 1 $ 15 $ 9
Product 2 20 4

a. What total monthly sales revenue is required to break even if the relative sales mix is 30 percent for Product 1 and 70 percent for Product 2? (Round your answer to the nearest dollar amount.)

b. What total monthly sales revenue is required to earn a monthly operating income of $16,000 if the relative sales mix is 20 percent for Product 1 and 80 percent for Product 2? (Round your answer to the nearest dollar amount.)
Answer #1

a) Weighted average contribution margin per unit = (9*30%+4*70%) = 5.5

Break even unit = 62000/5.5 = 11273 Unit

Product 1 = 11273*30% = 3382 Units

Product 2 = 11273*70% = 7891 Units

a) Weighted average contribution margin per unit = (9*20%+4*80%) = 5

Break even unit = 62000/5 = 12400 Unit

Product 1 = 12400*20% = 2480 Units

Product 2 = 12400*80% = 9920 Units